Loretta's age in 7 years will be three times what it was 9 years ago. How old is Loretta now?
Answer by stanbon(75887)

Loretta's age in 7 years will be three times what it was 9 years ago.
How old is Loretta now?
----------------
Equation:
(L+7) = 3(L-9)
---
L+7 = 3L - 27
2L = 34
Loretta is now 17 yrs old
===============================
Cheers,
Stan H.